When using the pooled variance t-test, if the sample size for the first group is 16 and the sample size for the second group is 25, how many degrees of freedom do we have when finding the critical values?
Answer #1

When we finding the critical value, the degree of freedom is given by

Given n1=16, n2= 25

Df = n1+ n2 - 2

= 16+25 - 2

Df = 41-2

Df =39